In a collection of nickels, dimes, and quarters the value of the collection is $12.85. If the number of nickels is 6 less than twice the number of dimes and the number of quarters is 5 more thsn twice the number of dimes,how many nickels are there?
Answer by mananth(16949)

In a collection of nickels, dimes, and quarters the value of the collection is $12.85. If the number of nickels is 6 less than twice the number of dimes and the number of quarters is 5 more thsn twice the number of dimes,how many nickels are there?
let dimes be x
value of dimes = 10x
nickels--------2x-6
value=5(2x-6)
Quarters2x+5
value=25(2x+5)
Total value=12.85 = 1285 cents
10x+5(2x-6)+25(2x+5)=1285
10x+10x-30+50x+125=1285
70x+95=1285
70x=1285-95
70x=1190
/70
x=17-dimes
nickels= 2x-6=>34-6=28
Quarter's = 2x+5 =>34+5=39
CHECK
39*25+28*5+17*10=